Nevey:

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? −x − 2y = −4 3x + y = 12 −2y − 4 2y − 4 2y + 4 −2y + 4

Nevey:

@BenLindquist @Shadow

Nevey:

@Pixel

dude:

Okay, first do you know how to isolate x?

Nevey:

i might, but i probably forgot. if u start it i will probably be like 'wait ik this!'

dude:

\(−x − 2y = −4\) What is the opposite of subtracting 2y?

Nevey:

sorry, internet. that would be adding 2y

dude:

Right so add 2y on both sides \(-x-2y=-4\)

Nevey:

-2y + 2y = 0 so that is cancled out leaving -x = -4.

RichardWatterson7:

−x−2y=−4 Step 1: Add 2y to both sides. −x−2y+2y=−4+2y −x=2y−4 Step 2: Divide both sides by -1. −x −1 = 2y−4 −1 x=−2y+4 Answer: x=−2y+4

Nevey:

-4 + 2y would by 2y.

dude:

Right on

Nevey:

be*

dude:

@RichardWatterson7 Please refrain from giving out direct answers.

RichardWatterson7:

💁👌

Nevey:

is there more to is? :/

dude:

\(-x = -4+2y\) We need to get rid of the - in front of the x What is the opposite of multiplying -1?

Nevey:

dividing -1

dude:

Right, so do it on both sides

Nevey:

could u say the problem again so i can see it? every time i try to copy it it shoots me back down here and i cant see it

dude:

\(−x=−4+2y\)

Nevey:

thx -1 divided by -x would cancel it out leaving it as just X

Nevey:

meaning it is now x = - 4 + 2y

Nevey:

@dude im stuck here lol

dude:

You were supposed to divide -1 on both sides So divide `-4+2y` by -1

Nevey:

-4 divided by -1 would by normal 4

dude:

Yes

Nevey:

do we do 2y divided by -1 or no? :/

dude:

Yes do 2y divided by -1

Nevey:

ok, that would be -2y

dude:

Correct, so now we have \(x=4-2y\)

dude:

Lets get the second equation given to us from the question \(3x + y = 12\)

dude:

Now substitute the newly found equation into the second equation, so you know how to do that?

Nevey:

no i donut :/

Nevey:

or i might lol

dude:

Try it

Nevey:

yea no i got nothin lol

dude:

Okay so we have the equations \(3\color{red}{x}+y=12\) and \(\color{red}{x}=4−2y\)

dude:

\(3\color{red}{x}+y=12\) \(3\boxed{\color{red}{4-2y}}+y=12\) Do you see how I did that?

Nevey:

i want to say that i do, but i dont :(

dude:

Let me try to color-code it  If you cannot still understand it, tell me which part confuses you

Nevey:

oh ok. no i got it now. what next?

dude:

Distribute the 3

Nevey:

3 x 4 is 12 and 3 x 2 is 6 + y would be 6y so it is now 12 - 6y + y = 12

dude:

Right on :P Now combine the y's \(12 - 6y + y = 12\)

Nevey:

12 - 6y + y = 12 y - y = 0 12 - 6 = 12 ( i think i did that right)

RichardWatterson7:

Nah u have to add well no subtract the y's then what do u get

dude:

Not quite You want to first combine the terms with the y

Nevey:

Shut up Richard. (those who have watched tommy boy will get it)

Nevey:

y = 1 right?

RichardWatterson7:

yea it does

Nevey:

ok thx Richard.

RichardWatterson7:

np math is my favorite subject sooo

Nevey:

yea i dont freakin get it

RichardWatterson7:

pretty much do 6y-y or 6y-1y and what do u get

Nevey:

5y

RichardWatterson7:

yea now what do u have

Nevey:

12 - 5y = 12

Nevey:

12 - 12 = 0 y = 5

RichardWatterson7:

No, subtract the 12 and what do u get

RichardWatterson7:

Do −5y+12−12=12−12 and what do u get